100cm3 of M sulphuric avid is need to an excess of solid sodium hydrogen trioxocarbonate (IV)


Calculates


1- The mass of sodium tetraoxosulphuric (VI) produce

1
Expert's answer
2021-12-08T01:44:02-0500

H2SO4 + 2NaHCO3 → Na2SO4 + 2H2O + 2CO2 (Balanced)


limiting reagent= H2SO4 because it gets used up.


Thus,


1 mole of H2SO4 → 1 mole of Na2SO4

1cm³ of H2SO4 → 1cm³ of Na2SO4

100cm³ of H2SO4 → 100cm³ of Na2SO4


molar mass of H2SO4 = 2 × 1 + 32 + 16 × 4= 98g/mol

molar mass of Na2SO4= 2 × 23 + 32 + 16 × 4= 142g/mol


98g of H2SO4 → 142g of Na2SO4


mass= no of moles × molar mass


mass of Na2SO4 produced= 100 × 142= 14200g.